OpenStudy (anonymous):

solve 125x^3 + 343 = 0

OpenStudy (anonymous):

\[125 x^3+343=(5 x+7) \left(25 x^2-35 x+49\right) \]

OpenStudy (anonymous):

omg thank you moneybird

OpenStudy (anonymous):

now you are being nice :D

OpenStudy (anonymous):

i wasnt ever being mean, you were just being rude

OpenStudy (anonymous):

125x^3 + 343 = 0 125x^3 = -343 taking cube root on both sides 5x = -7 => x= -7/5

OpenStudy (anonymous):

i guess i just dont know how to take a cube root, that is my problem

OpenStudy (anonymous):

i was joking :3

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

im just having a rough night

OpenStudy (anonymous):

Do you need any more help?

OpenStudy (anonymous):

maybe, idk yet

OpenStudy (anonymous):

you dont know to take cube root ??????

OpenStudy (anonymous):

no! i dont! okay!

OpenStudy (anonymous):

She's very sensitive

OpenStudy (anonymous):

haha, well you dont know what i went through tonight

OpenStudy (anonymous):

What did you do?

OpenStudy (anonymous):

its not what i did but what people did to me, i dont really wanna talk about it

OpenStudy (anonymous):

whos very sensitive @moneybird

OpenStudy (anonymous):

not you

OpenStudy (anonymous):

!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

aww poor kamc23

OpenStudy (anonymous):

thanks :)